OpenStudy (anonymous):

Match the information about interest earned with the principal amount invested. Remember the equation for calculating simple interest is I equals p times r times t. Match Term Definition eighty−one dollars over 3 years at an interest rate of three percent A) p equals one thousand one hundred dollars eighty dollars over 2 years at an interest rate of five percent B) p equals eight hundred dollars ninety−six dollars over 2 years at an interest rate of four percent C) p equals nine hundred dollars ninety−nine dollars over 3 years at an interest rate of three percent D) p equals one tho

OpenStudy (anonymous):

can i just get the answere

OpenStudy (texaschic101):

I = PRT In these problems, you are looking for P....so lets change the formula to : P = I/RT =============== I = 81 T = 3 R = 3% = 0.03 (always change rate to decimal by dividing by 100) now sub P = 81/(0.03)(3) P = 81 / .09 P = 900 -- answer C ....thats how you would do them rest of them... P = I/RT P = 80/(0.05)(2) P = 80/0.1 P = 800 ======= P = I/RT P = 96 / (.04)(2) P = 96 / 0.08 P = 1200 ====== P = I/RT P = 99/(.03)(3) P = 99 / .09 P = 1100